# Adacc: Adaptive Compression and Activation Checkpointing for LLM Memory Management

### v1

@@ Abstract.
Training large language models often employs recomputation to alleviate memory pressure, which can introduce up to 30% overhead in real-world scenarios. In this paper, we propose Adacc, a novel memory management framework that combines adaptive compression and activation checkpointing to reduce GPU memory footprint. It comprises three modules: (1) We design layer-specific compression algorithms that account for outliers in LLM tensors, instead of directly quantizing floats from FP16 to INT4, to ensure model accuracy. (2) We propose an optimal scheduling policy that employs MILP to determine the best memory optimization for each tensor. (3) To accommodate changes in training tensors, we introduce an adaptive policy evolution mechanism that adjusts the policy during training to enhance throughput. Experimental results show that Adacc can accelerate the LLM training by 1.01 to 1.37 compared to state-of-the-art frameworks, while maintaining comparable model accuracy to the Baseline.
@@ 1. Introduction
Limitations of existing works. The existing recomputation methods can be placed into two categories. First, coarse-granularity recomputation selects specific tensors for recomputation based on predefined rules, following an all-or-nothing pattern. For instance, Megatron libraries offer a full recomputation option that caches only the inputs to transformer layers as checkpoints, discarding other activations and recomputing them during backward propagation. However, this approach introduces significant recomputation overhead. Specifically, we profile Megatron-LM using a 345M parameter GPT model with an 8 batch size on 4 Tesla V100 GPUs. Our results indicate that full recomputation accounts for over 28% of the total training time. This inefficiency arises because the rule-based policy does not account for the varying memory demands and computational costs of different model layers, as described in §2.2
Second, model-adaptive recomputation methods (Chen et al., 2016, 2024) have been proposed to employ search algorithms to identify the optimal policy based on the characteristics of deep learning models and the capabilities of the devices. For example, Checkmate uses MILP to find a suitable recomputation policy through offline searching. However, due to the vast search space, it is unable to find the optimal solution for the large language model.
| Tensor | Size (MB) | Recmp. time (ms) | Compression time (ms) | Compression overhead (MB) | Recmp. Comp. |
@@
*Recmp. time: the time recomputed from the previous layer to this layer. *As recomputation frees tensors from the GPU, the space overhead is zero.
New observations. Data compression techniques have been proposed to reduce data size. Surprisingly, advancements in GPU parallel computing and algorithm optimization have significantly improved computation efficiency, making data compression feasible for AI training. We explore compression opportunities by selecting activation tensors from the GPT model and presenting the computational time and memory costs for recomputation and the proposed compression methods (in §3.2). As shown in Table 1, We identify four tensors, each suited to different memory optimization techniques. Specifically, we define the optimization bandwidths, and , to demonstrate the efficiency of each technique. The bandwidth is calculated as . For example, the and for T1 is 266 and 194, respectively. It is better to optimize T1 by recomputation. In contrast, compression is more efficient for T2. Therefore, different memory optimization strategies should be applied to tensors to maximize model training throughput rather than only considering the recomputation as implemented in mainstream frameworks (NVIDIA, 2024b).
Our work. This finding motivates the proposal of a new adaptive memory management framework, Adacc, that integrates compression and recomputation to maximize training throughput and prevent GPU out-of-memory issues. However, Adacc faces three challenges when introducing compression into the training framework. First, directly applying compression into framework can lead to greater loss of model accuracy. To address this, Adacc designs four layer-specific compression algorithms that account for data outliers, offering fast, memory-efficient solutions with minimal impact on training accuracy. Second, how to find the optimal combined policy for compression and recomputation within an acceptable time? To solve it, Adacc uses a custom mixed-integer linear programming (MILP) algorithm to achieve a globally optimized tensor policy for LLM models, significantly reducing the search space by leveraging the repeated training structures in LLMs. Third, the dynamic variation of data outliers during training affects the model’s compression rate, rendering the current policy suboptimal. Therefore, Adacc adaptively adjusts the model policy efficiently in response to data changes to enhance training efficiency.
We evaluate Adacc on a wide range of benchmarks and downstream tasks across diverse scenarios, achieving significant performance gains without compromising model accuracy.
@@ 2.1. Large language model training
However, GPUs’ limited memory capacity restricts large model training, as both model states and activations (feature maps) consume memory. Model states consist of parameters, gradients, and optimizer states, such as momentum and variance in algorithms like Adam. Activations are intermediate data generated during training that need to be stored on the GPU, with memory consumption increasing as batch size grows. Users often use large training batches to maximize GPU utilization, which exacerbates memory bottleneck issue (i.e., memory wall). For example, in GPT model training, 66% of memory is consumed by stashed immediately activations on 4 V100 GPUs with a batch size of 4. This number grows to 76% for 8 batch size. Therefore, optimizing activations is essential for efficient LLM training.
@@ 2.2. Existing works and their shortcomings
Several works have explored optimizations for GPU memory wall issue. Prior wroks use activation checkpointing (recomputation) to extend the limited capacity of GPU memory (Chen et al., 2016; Jain et al., 2019; Chen et al., 2024). However, we find not all activations are suited for recomputation. Small activations that are costly to recompute should remain in GPU memory or be optimized with other techniques, as tensors within the red circle shown in Figure 1. Some works propose lossless and lossy encodings to reduce activation memory (Jain et al., 2018; Rhu et al., 2018), but these target traditional CV models, whereas Adacc focuses on large language models. Some methods swap activations from GPU to CPU to save memory (Rhu et al., 2016; Chen et al., 2022), but they suffer performance loss due to limited PCIe bandwidth and require intrusive framework modifications, making them impractical. Additionally, intra- and inter-operator parallelism methods split models and distribute partitions across GPUs to reduce memory footprints (Rajbhandari et al., 2020; Shoeybi et al., 2019; Huang et al., 2019), complementing Adacc.
@@ 3.1. Overview
3.2. Layer-specific Compression
Traditional compressions. A straightforward approach is to compress activations from FP16 to INT4 using per-token quantizaiton (as shown in Figure 3(b)(i)). However, our experience shows that this naive approach can reduce model accuracy by up to 39% in GPT models or even leads to gradient overflow in training (as shown in §4.3). To solve this challenge, we design several specific compression algorithm for different LLM layers.
Outlier-separated activation compression. As shown in Figure 3(a), there are many outliers in LLM tensors. Existing methods quantize outliers with other values, resulting in significant outliers accuracy loss and greater model degradation. To address this, a naive method groups activations by channel and quantize outliers separately from the rest of the values, as shown in Figure 3(b)(ii). However, since the data primitively is stored along token dimension (row), data adjacent on the channel dimension (column) is not contiguous in memory. Since the compressor requires contiguous data in memory, we need to transpose and align the tensor before compression. However, performing these operations can be time-consuming. Therefore, we design an outlier-separated method to isolate the outlier from normal value, and quantize the normal layer based on their original alignment, as shown in Figure 3(b)(iii). We use Z-Score method to extract outlier channels from activations, and the Z-Score for each channel is calculated using formulation:
Layer-specific compression. We apply different compression scheme to different activations based on their features.
For Linear, LayerNorm, and GELU layers with many outliers, we apply an outlier-separated activation compression scheme to minimize errors.
3.3. Determining Cost-Effectiveness of Tensor Optimization
@@
As analysis in §2.2, selecting an appropriate memory optimization policy for different tensors is crucial for achieving optimal training efficiency. Each layer can opt for either layer-specific compression or recomputation. To find the policy, existing approaches build cost models for all operators, resulting in a large search space, which is impractical for large language models (Jain et al., 2019). Surprisingly, LLMs are composed of repeated identical structures, such as transformer blocks in GPT, with consistent memory and computation requirements, as well as a similar outlier rate. As a result, the locally optimal recomputation policy for a single layer can be applied to other identical layers without searching the global space.
@@
MILP cost model. We use a mixed-integer linear program (MILP) to formulate and describe the problem. Here, we illustrate the design with a transformer-based model. However, our approach also applies to other large deep learning models.
@@ 3.4. Adaptive Policy Evolution
As shown in Figure 4, the outliers in tensors vary across different training stages, causing the compression ratio to change dynamically during the training process. Therefore, we must track outliers changes to update memory optimization policy and maximize training throughput. While, tracking outliers requires traversing the entire tensor, making it time-consuming. This introduces a challenge: high-frequency tracking incurs more overhead but improves policy efficiency, while low-frequency tracking minimizes overhead but leads to suboptimal training throughput. Therefore, how to design an optimal tracking frequency is important.
@@
Policy evolution mechanism. During training, at each tracking iteration, the framework re-evaluates outliers in the Linear, LayerNorm, and GELU layers and inputs the updated tensor information (e.g., ) into the MILP cost model. If the memory optimization policy changes, the framework adjusts it and continues training with the updated policy.
@@ 4.2. General Result
| Model | Baseline | Full Recomp. | Quantization | Adacc |
| GPT-117M | 57 | 136 | 98 | 136 |
| GPT-345M | 8 | 61 | 19 | 61 |
| GPT-4.7B | 1 | 4 | 3 | 4 |

### v2

# Adacc: An Adaptive Framework Unifying Compression and Activation Recomputation for LLM Training

@@ Abstract.
Training large language models (LLMs) is often constrained by GPU memory limitations. To alleviate memory pressure, activation recomputation and data compression have been proposed as two major strategies. However, both approaches have limitations: recomputation introduces significant training overhead, while compression can lead to accuracy degradation and computational inefficiency when applied naively.
In this paper, we propose Adacc, the first adaptive memory optimization framework that unifies activation recomputation and data compression to improve training efficiency for LLMs while preserving model accuracy. Unlike existing methods that apply static, rule-based strategies or rely solely on one technique, Adacc makes fine-grained, tensor-level decisions, dynamically selecting between recomputation, retention, and compression based on tensor characteristics and runtime hardware constraints. Adacc tackles three key challenges: (1) it introduces layer-specific compression algorithms that mitigate accuracy loss by accounting for outliers in LLM activations; (2) it employs a MILP-based scheduling policy to globally optimize memory strategies across layers; and (3) it integrates an adaptive policy evolution mechanism to update strategies during training in response to changing data distributions. Experimental results show that Adacc improves training throughput by 1.01 to 1.37 compared to state-of-the-art frameworks, while maintaining accuracy comparable to the baseline.
@@ 1. Introduction
Existing recomputation methods often incur substantial training overhead due to their rigid, rule-based design (Chen et al., 2024). These approaches typically select specific tensors for recomputation based on predefined heuristics, following an all-or-nothing pattern (Deepspeed, 2023). For example, Megatron-LM provides a full recomputation mode, which retains only the inputs to transformer layers as checkpoints and discards all other activations, requiring them to be recomputed during the backward pass. Specifically, we profile Megatron-LM using a 345M parameter GPT model with an 8 batch size on 4 Tesla V100 GPUs. Our results indicate that full recomputation accounts for over 28% of the total training time. This inefficiency arises because the rule-based policy does not account for the varying memory demands and computational costs of different model layers.
| Tensor | Size (MB) | Recmp. time (ms) | Compression time (ms) | Size after compression (MB) | Recmp. Comp. |
@@
*Recmp. time: the time required to recompute from the previous layer.
In addition to recomputation, data compression has emerged as another mainstream approach for reducing memory usage during model training. A common strategy is to apply quantization to compress activations after the forward pass and decompress them before the backward pass, thereby reducing memory consumption without modifying the computational graph (Chen et al., 2021). However, even simple quantization schemes introduce non-negligible overhead due to the additional computation required for both quantization and dequantization, which can negatively impact overall training efficiency.
We find that different activation tensors are best optimized using different memory strategies—some favor recomputation, while others benefit more from compression. This variation arises from intrinsic differences in tensor characteristics, such as their size and computational cost. To validate this, we select several representative tensors from a GPT model and compare their memory and time trade-offs under both optimization methods.
As shown in Table 1, we evaluate four representative tensors, each favoring a different memory optimization strategy. To quantify the trade-offs, we define the optimization bandwidths, and , as the saved memory per unit time, calculated as , where denotes the actual memory footprint under each strategy (e.g., zero for recomputation and compressed size for compression). This metric reflects the memory-efficiency of each technique in terms of throughput. For instance, T1 yields a of 266 and a of 194, indicating that recomputation is more favorable. In contrast, T2 exhibits higher efficiency under compression, making it the better choice for that tensor.
These results demonstrate that applying a one-size-fits-all policy—whether solely relying on recomputation or compression—is suboptimal. Instead, tensor-aware hybrid strategies that dynamically adapt to the properties of individual tensors are essential for maximizing training throughput, challenging the static, rule-based approaches employed in mainstream frameworks such as Megatron (NVIDIA, 2024b).
This insight motivates an adaptive framework combining compression and recomputation to maximize training efficiency and mitigate GPU memory limitations. However, it faces three key challenges. First, directly applying compression may lead to significant accuracy degradation. Second, it is non-trivial to determine the optimal combination of compression and recomputation strategies within a reasonable time budget. Third, the effectiveness of compression depends on the statistical properties of activations, such as the presence of outliers, which can vary dynamically during training. As a result, a fixed compression policy may become suboptimal as training progresses.
To address these challenges, we propose Adacc, which first introduces four layer-specific compression algorithms tailored to different types of activations; notably, one of them explicitly handles data outliers to improve compression robustness, while others adopt lightweight strategies for efficiency. It then employs a custom mixed-integer linear programming (MILP) algorithm to derive globally optimized tensor-level policies, significantly reducing the search space by leveraging the repetitive structure of LLMs. Finally, Adacc adaptively updates its optimization policy during training to accommodate dynamic data characteristics and sustain high throughput.
We conduct extensive evaluations of Adacc on a broad range of benchmarks and downstream tasks, showing that it consistently improves training performance without compromising model accuracy.
@@ 2.1. Large language model training
However, GPUs’ limited memory capacity restricts large model training, as both model states and activations (feature maps) consume memory (Rajbhandari et al., 2020). Model states consist of parameters, gradients, and optimizer states, such as momentum and variance in algorithms like Adam (Kingma and Ba, 2014). Activations are intermediate data generated during training that need to be stored on the GPU, with memory consumption increasing as batch size grows. Users often use large training batches to maximize GPU utilization, which exacerbates memory bottleneck issue (i.e., memory wall). For example, in GPT model training, 66% of memory is consumed by stashed immediately activations on 4 V100 GPUs with a batch size of 4. This number grows to 76% for 8 batch size. Therefore, optimizing activations is essential for efficient LLM training.
@@ 2.2. Existing works and their shortcomings
Recomputation. Several studies have addressed the GPU memory wall issue by exploring various optimization techniques aimed at alleviating the constraints imposed by limited GPU memory. One common approach involves activation checkpointing (recomputation), which temporarily discards intermediate activations and re-computes them during the backward pass to save memory (Chen et al., 2016; Jain et al., 2019; Chen et al., 2024). While this method is effective in extending GPU memory capacity, we observe that not all activations are suitable for recomputation. Specifically, small activations that are costly to recompute should either remain in GPU memory or be optimized using alternative techniques, as highlighted by the tensors within the red circle in Figure 1. These activations present unique challenges that are often overlooked in traditional checkpointing strategies.
Compression. Further attempts to address memory limitations have involved both lossless and lossy encoding techniques, which compress activation data to reduce memory consumption (Jain et al., 2018; Rhu et al., 2018). However, these approaches are typically designed for traditional computer vision (CV) models, and they do not directly cater to the unique memory demands of large language models (LLMs), which have more complex and larger activation patterns. In contrast, Adacc is specifically tailored to optimize memory usage in the context of LLMs, addressing these specific challenges with greater efficiency.
Data swapping. Another approach involves swapping activations between GPU and CPU memory, which can provide additional memory space on the GPU (Rhu et al., 2016; Chen et al., 2022). However, this technique introduces significant performance overhead due to the limited PCIe bandwidth, and often requires intrusive modifications to the underlying framework, making it impractical for real-time or large-scale training scenarios (Liao et al., 2024).
Parallelism policy.In addition, methods that leverage intra- and inter-operator parallelism have been proposed to distribute model partitions across multiple GPUs, thereby reducing the memory footprint on each individual GPU (Rajbhandari et al., 2020; Shoeybi et al., 2019; Huang et al., 2019). These methods complement Adacc, providing an alternative approach to memory optimization by spreading the computational load. However, they still face challenges in terms of coordination and communication overhead between GPUs.
In summary, while various methods have been proposed to address the GPU memory wall issue, each comes with trade-offs in terms of computational cost, practicality, and the specific requirements of large-scale models. Adacc distinguishes itself by offering a more tailored solution for the memory optimization needs of large language models.
@@ 3.1. Overview
3.2. Layer-specific Compression
Traditional compressions. A straightforward approach is to compress activations from FP16 to INT4 using per-token quantizaiton (as shown in Figure 3(b)(i)). However, our experience shows that this naive approach can reduce model accuracy by up to 39% in GPT models or even leads to gradient overflow in training (as shown in evaluation section). To solve this challenge, we design several specific compression algorithm for different LLM layers.
Outlier-separated activation compression. As illustrated in Figure 3(a), activations in LLMs often exhibit a heavy-tailed distribution, with a small number of extreme outlier values. These outliers can carry important semantic or gradient information, particularly in attention and normalization layers. However, existing quantization methods typically apply uniform quantization across the entire tensor without isolating outliers, forcing them to share quantization levels with densely populated regions of the value distribution. This results in high quantization error for the outliers, leading to loss of critical information and significant model degradation.
To address this, a naive method groups activations by channel and quantizes outliers separately from the rest of the values, as shown in Figure 3(b)(ii). However, since the data is primitively stored along the token dimension (row), data adjacent on the channel dimension (column) is not contiguous in memory. Since the compressor requires contiguous data in memory, we need to transpose and align the tensor before compression. Unfortunately, performing these operations can be extremely resource-intensive and significantly slow down the process. Through our observations, we noticed that isolating outliers from normal values significantly improves efficiency. This is because, by separating the outliers, we can avoid the need for tensor transposition and alignment, which are computationally expensive. Without the need to rearrange the data in memory, the normal values can be quantized directly while maintaining their original alignment, thus reducing the overhead.
Based on this insight, we propose an outlier-separated method that isolates the outliers from the normal values and quantizes the normal values without the need for additional transformations, as shown in Figure 3(b)(iii). This approach streamlines the process and avoids the performance bottleneck caused by tensor manipulation.
To efficiently isolate outlier channels from activations, we designed a Z-Score method. The Z-Score is a statistical measure that quantifies how far a particular data point (or in this case, an activation channel) deviates from the mean in terms of standard deviations. This method is crucial for distinguishing extreme values from the normal distribution of the data. The Z-Score for each channel is calculated using the following formula:
Layer-specific compression. We apply different compression schemes to activations according to their statistical and computational properties. Importantly, the targeted layer types are core components in most large language models, including GPT, BERT, T5, and LLaMA, which makes our method architecture-agnostic and widely applicable.
For Linear, LayerNorm, and layers with many outliers, we apply an outlier-separated activation compression scheme to minimize errors.
3.3. Determining Cost-Effectiveness of Tensor Optimization
@@
As analyzed in the motivation section, selecting an appropriate memory optimization policy for different tensors is crucial for achieving optimal training efficiency. Each layer can opt for either layer-specific compression or recomputation. To determine the optimal policy, existing approaches often construct cost models over the full computation graph (Jain et al., 2019), resulting in a large combinatorial search space that is impractical for large language models (LLMs).
Challenges. Designing this optimization problem presents several challenges. First, recomputation decisions must respect data dependencies: a layer can only be recomputed if all its upstream layers are either stored or also marked for recomputation, introducing nontrivial inter-layer constraints. Second, compression introduces heterogeneous trade-offs, different layers employ distinct compression schemes with varying memory savings and computational overheads, which must be precisely modeled to ensure accurate cost estimation. Together, these factors not only complicate the problem formulation but also significantly enlarge the solution space, making the optimization process more challenging.
Fortunately, LLMs consist of repeated, structurally identical blocks—such as transformer layers in GPT, with consistent memory and data patterns. This regularity allows us to derive a locally optimal policy for one block and apply it across others, significantly reducing the search space without compromising policy quality.
@@
MILP cost model. We formulate the optimization as a mixed-integer linear program (MILP), leveraging the repeated structural patterns commonly found in LLMs to reduce the problem complexity. For illustration, we describe the design using a transformer-based model; however, the approach generalizes to other large-scale deep learning models with similar modular architectures.
@@
| Model | Baseline | Full Recomp. | Quantization | Adacc |
| GPT-117M | 57 | 136 | 98 | 136 |
| GPT-345M | 8 | 61 | 19 | 61 |
| GPT-4.7B | 1 | 4 | 3 | 4 |
@@ 3.4. Adaptive Policy Evolution
However, monitoring outlier patterns incurs non-negligible overhead due to full-tensor traversal and system queries. For example, when tracking outliers across large neural network layers, the process may require scanning through millions of tensor elements and frequently querying the system for updates. This results in significant computational and memory costs. Consequently, there is a trade-off: frequent tracking improves policy accuracy by capturing real-time changes in the model’s behavior, but at the expense of increasing the overall computational burden and slowing down training. On the other hand, infrequent tracking reduces the overhead and allows for faster processing, but it comes at the risk of missing critical shifts in the model’s performance, leading to potentially suboptimal decisions. For instance, in large language models, this could mean missing key activations that impact model convergence or accuracy. Thus, determining an appropriate tracking frequency is key to balancing accuracy and efficiency, ensuring that the system remains responsive while avoiding excessive computational costs.
@@
Policy evolution mechanism. During training, at each tracking iteration, the framework re-evaluates outliers in the Linear, LayerNorm, and GELU layers and inputs the updated tensor information (e.g., ) into the MILP cost model. Additionally, the framework allows users to customize which layers they want to re-evaluate, providing flexibility in targeting specific layers for dynamic adjustments during the training process. If the memory optimization policy changes, the framework adjusts it and continues training with the updated policy.